### v1

# Explainability of AI Uncertainty: Application to Multiple Sclerosis Lesion Segmentation on MRI

Abstract
Trustworthy artificial intelligence (AI) is essential in healthcare, particularly for high–stakes tasks like medical image segmentation. Explainable AI and uncertainty quantification significantly enhance AI reliability by addressing key attributes such as robustness, usability, and explainability. Despite extensive technical advances in uncertainty quantification for medical imaging, understanding the clinical informativeness and interpretability of uncertainty remains limited. This study introduces a novel framework to explain the potential sources of predictive uncertainty, specifically in cortical lesion segmentation in multiple sclerosis using deep ensembles. The proposed analysis shifts the focus from the uncertainty–error relationship towards relevant medical and engineering factors. Our findings reveal that instance–wise uncertainty is strongly related to lesion size, shape, and cortical involvement. Expert rater feedback confirms that similar factors impede annotator confidence. Evaluations conducted on two datasets (206 patients, almost 2000 lesions) under both in–domain and distribution–shift conditions highlight the utility of the framework in different scenarios.
keywords:
Uncertainty quantification , Instance–wise uncertainty , Explained uncertainty , Explainable AI , Lesion segmentation , Multiple sclerosis , Magnetic resonance imaging
journal:
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Addresses clinical explainability gaps in medical AI uncertainty methods.
Novel framework explains AI uncertainty in MS lesion segmentation.
Lesion–wise deep ensemble uncertainty is explained through lesion and patient factors.
Explanations correspond to the factors impending expert rater confidence.
Extensive validation across two large cortical lesion MRI datasets.
1 Introduction
Trustworthy artificial intelligence (AI) refers to AI systems designed and implemented to meet essential ethical and performance standards, ensuring safe and beneficial integration into society [1, 2]. Ensuring AI trustworthiness is crucial in high-risk fields such as healthcare, where AI-driven decisions directly impact patient outcomes [3, 4]. Regulatory efforts, including the European Union’s AI Act [5], emphasize transparency, accountability, and reliability, mandating clear documentation and interpretability of AI systems to safeguard patient welfare and rights. Recent international guidelines for AI in healthcare, FUTURE-AI [6], have highlighted fairness, universality, traceability, usability, robustness, and explainability as key characteristics needed in the development and deployment of trustworthy AI in healthcare.
Explainable AI (XAI) methods [7] and uncertainty quantification (UQ) [8] play a critical role in trustworthy AI by directly addressing several key attributes outlined by international guidelines and regulatory bodies. XAI methods primarily improve traceability, usability, and explainability by providing insights into how models arrive at specific predictions [7, 6]. UQ significantly contributes to robustness, usability, and explainability by quantifying the reliability of model predictions, highlighting areas of potential model failure, and assisting users in interpreting ambiguous outcomes [8, 6]. Together, UQ and XAI methods facilitate informed interactions between humans and AI, promoting trust and supporting robust clinical decision-making [9].
In this context, UQ has become increasingly important in medical image analysis, particularly in tasks involving deep learning (DL) based segmentation of pathological features [10, 4, 11]. UQ assists clinicians by identifying ambiguous or unreliable predictions, guiding them to areas needing manual review, thus improving clinical confidence and decision quality. However, effectively evaluating uncertainty is inherently challenging due to the absence of definitive ground truth and the diversity of methods used to quantify uncertainty [8]. Although substantial technical progress has been achieved, the clinical impact of UQ remains underinvestigated [12, 13]. Common evaluation strategies assess the relationship of quantified uncertainty with error but do not reveal whether uncertainty is associated with clinically meaningful factors (e.g., lesion characteristics, image quality) or if it aligns with the expert definition of confidence [4]. To bridge this gap, UQ needs to be evaluated not only for technical performance but also for clinical informativeness and capacity to improve human-AI collaboration [14, 9].
In this work, we explore the trustworthy behaviour of DL-based neuroimage analysis to support the assessment of multiple sclerosis (MS) [15, 16]. These patients suffer from a chronic neurodegenerative disease, subsequently evaluated with the use of magnetic resonance imaging (MRI) biomarkers that can be automatically assessed with ad-hoc AI systems. The availability of recommendations and guidelines makes the automated segmentation of MS lesions an ideal scenario to further explore the expert explainability [16, 15, 17]. Focal lesion segmentation in MS serves as an example application to study UQ and explainability, given its inherent complexity and multiple sources of uncertainty [18]. Model uncertainty arises from limited training data and distributional shifts, whereas the irreducible data uncertainty is driven by the data noise and the ambiguity of the definition of the pathology [8]. This complexity intensifies for recently incorporated biomarkers in the diagnostic criteria for MS, such as cortical lesions (CLs), paramagnetic rim lesions, and central vein sign, where sources of uncertainty are magnified due to their small size, subtle contrast differences, and complexity in segmentation [19]. Thus, AI models that can reliably quantify these advanced biomarkers and integrate them safely into the patient assessment workflow are needed to diagnose, treat, and understand the disease course.
1.1 Contributions
In this study, we propose a novel framework to explain AI uncertainty beyond its relationship with prediction errors, specifically within the context of CL segmentation in MS neuroimaging. The proposed framework is used to evaluate the statistical dependencies between deep ensemble lesion-wise uncertainty and both lesion-specific and patient-specific features. Our framework aims to: (i) explain model behavior, (ii) perform uncertainty sanity checks, and (iii) differentiate between informative uncertainty that supports decision-making and misleading uncertainty that could adversely influence clinical interpretations. Through expert rater feedback, we investigated factors influencing annotator confidence and explored their intersection with high-uncertainty factors. We conducted the analysis across two comprehensive datasets, containing 206 patients and approximately 2000 CLs, evaluated both in-domain and under distribution shifts, extending our previous work in an important way [20].
2 Related work
2.1 Lesion segmentation in MS
MS is a chronic inflammatory and neurodegenerative disease with a prevalence of 2 million worldwide, characterized by focal, inflammatory demyelination in the central nervous system [15, 16]. MRI plays a key role in MS diagnosis, prognosis, and treatment planning [17]. White matter lesions (WML) detectable on MRI are a hallmark of MS, included in the diagnostic criteria and follow-up pipelines [16]. DL models have been widely explored to automate the tedious and time-consuming annotation process of WML annotation on MRI [21, 22, 23]. Several studies  [23, 24] have highlighted the utility of U-Net-like architectures  [25, 26], dominated by nnU-Net [27, 28] in recent years. CLs have emerged as a promising MRI biomarker that supports the differential diagnosis of MS and aids in predicting disease progression [16, 29, 19, 30, 31]. The CL segmentation on MRI presents a greater challenge both for radiologists and machine learning models, given their small size (as small as a few voxels or milliliters), ambiguity in CL definition and confusion with the WML juxtacortical lesions (see Figure 7). Due to the lack of publicly available data and the complexity of the task, the automation of the CL segmentation task has not been widely explored. Some articles have proposed joint segmentation of CLs and WMLs [32, 33], for which the performance can be dominated by the more prevalent and visible WMLs. CL segmentation was explored in the context of ultra-high field MRI [34] and comparing the visibility of CL on several advanced MRI modalities [35].
2.2 Uncertainty quantification for medical image segmentation
UQ is widely adopted within the general field of medical image segmentation [10, 4, 11]. A range of methodological approaches have been proposed, including Bayesian methods, ensemble-based methods, conformal prediction and others (such as test-time augmentations, heteroscedastic models, and uncertainty prediction) [10, 4, 11]. Methods such as Monte Carlo dropout [36] and deep ensembles [37] dominate the landscape, with ensembles providing superior uncertainty calibration, albeit at a higher computational cost [4]. To mitigate this, single-inference approaches were explored, including temperature-scaled softmax probability calibration, evidential DL-based on Dempster–Shafer theory, and learned uncertainty frameworks, predicting uncertainty as one of the outputs [4]. In parallel, generative models—such as probabilistic U-Nets [38], PHISeg [39], or diffusion-based architectures [40] — were employed to sample diverse plausible segmentation masks, enabling fine-grained modeling of aleatoric uncertainty and inter-rater variability, although often at a higher computational cost. Test-time augmentation was proposed within the medical imaging field as a model-agnostic way to quantify aleatoric uncertainty [41]. Several approaches go beyond per-voxel uncertainty and propose to quantify uncertainty per region of interest (ROI), e.g., tumor [42] or whole brain [43, 44]. Commonly, these uncertainty measures average the voxel-wise uncertainty within a ROI [45, 42, 43] or quantify the discrepancies in the binary ROI prediction across different samples [42, 44].
For the task of WML MS segmentation trustworthiness assessment using UQ was previously explored using different techniques, including label-flip uncertainty prediction method [46, 47], Monte Carlo dropout [45, 48], batch ensembles [47] and deep ensembles [48, 44]. Several studies proposed to quantify uncertainty associated with each predicted lesion region to aid in the false positive lesion discovery[45, 49] or segmentation quality prediction[50]. Further usage of uncertainty varies, including trustworthiness assessment, quality control, data correction, active learning, and out-of-distribution detection [10, 4, 11].
2.3 Evaluating uncertainty quantification
UQ evaluation is an active area of research [51]. Assessing the relationship between uncertainty and error is a common practice, however, it does not unravel all the information brought by the uncertainty values [52]. A common evaluation pipeline relies on the assessment of uncertainty calibration (e.g., expected calibration error) [53] or correlation with error [42], robustness-uncertainty measures (e.g., error retention curves) [54, 55, 44]. Some works compare uncertainty against the inter-rater disagreement for segmentation tasks [38, 39]. Qualitative visual assessment of the uncertainty maps serves as a sanity check and enhances the interpretability of the uncertainty values. For instance, high uncertainty values are known to appear at the borders of ROIs [39, 46, 55, 44].
Clinical evaluation of uncertainty is limited to the perceptual assessment rather than clinical informativeness analysis [12, 13]. A pioneering study  [12] on the XAI and uncertainty perception in the clinics found that trust scores were the most positively perceived explanation type among pathologists, as they provided intuitive and immediately actionable confidence information. A more recent work  [13] studies the visualization scenarios of uncertainty maps in the clinics. This work reported that clinicians preferred binary structure-wise interpretable uncertainty visualizations, paired with voxel-level detail when needed.
Several studies have proposed explainability approaches to unravel the information behind the uncertainty values. One study on diabetic retinopathy images classification [56] analyzed the causes of uncertainty by examining the predictive variance across Monte Carlo dropout samples and visualizing its behavior to clinical evaluation or unfamiliar inputs, offering qualitative explanations of model uncertainty. Another study [57] used t-SNE to explain uncertainty in Bayesian CNN-based classification of breast histopathology images, showing how high-uncertainty inputs clustered in latent space and could be flagged for human review, while low-uncertainty cases exhibited significantly improved classification accuracy. To the best of our knowledge, the explainability of uncertainty has not been explored for medical image segmentation.
3 Materials and methods
3.1 Data
We used two datasets acquired at the i) University Hospital Basel [58] and ii) Lausanne University Hospital [59] in Switzerland. In total, we included data from 206 patients diagnosed with MS according to McDonald criteria [16], having different clinical phenotypes: relapsing-remitting (140), primary-progressive (22), and secondary-progressive (44). The dataset comprises 3D magnetization-prepared 2 rapid gradient-echo (MP2RAGE) images [60] and manual CL annotations. In University Hospital Basel, the imaging was performed using a standardized acquisition protocol on a 3 Tesla whole-body MR system (MAGNETOM Prisma, Siemens Healthcare) with a 64-channel head and neck coil. In Lausanne University Hospital, a 3 Tesla whole-body MR system was used (MAGNETOM Trio, Siemens Healthcare) with a 32-channel head and neck coil.
In both medical centers, annotations were obtained by the consensus: in University Hospital Basel, by a neurologist and a medical doctor with 5 and 5 years of experience in MRI; in Lausanne University Hospital, by a neurologist and a radiologist with 11 and 7 years of experience in MRI. The identified CLs encompassed both intracortical lesions (confined within the gray matter (GM)) and leukocortical lesions (involving both the cortical GM and the adjacent white matter (WM)).
The data from the University Hospital (n=163) was split into training, validation, and test sets with 109:13:41 subjects respectively, corresponding to 857:69:301 CLs. The data from Lausanne University Hospital was solely used for the out-of-distribution evaluation and contains 758 CLs. Further in the manuscript, the in-domain and out-of-domain test sets will be referred to as Test-in and Test-out. The distibutional shift between Test-in and Test-out is due to the change of the medical center, MRI scanner model, acquisition protocol, annotators, annotation protocols, and lesion intensity distributions. Examples of CLs are shown in Figure 1. The summary of the data information is given in Table 1.
Figure 1: Examples of cortical and white matter lesions and their subtypes on MP2RAGE MRI scan: intracortical, leukocortical, juxtacortical, deep white matter lesions. Lesions appear as hypointense regions within GM and WM. The examples illustrate a possible similarity between leukocortical and juxtacortical, contributing to the confusion between cortical and white matter lesions.
Table 1: Data description. RR - relapsing-remitting, PP - primary-progressive, SP - secondary-progressive; Q2 - median, IQR - interquartile range.
Dataset parameters In-domain Out-of-domain
Source University Hospital Basel Lausanne University Hospital
Disease phenotypes RR (97), PP (22), SP (44) RR
Number of MS patients 163 43
Total number of CLs 1227 758
M:F ratio 0.68 0.72
Scanner Magnetom Prisma, Siemens Healthineers Magnetom Trio, Siemens Healthineers
TR, TE, TI1, TI2, ms 5000, 2.98, 700, 2500 5000, 2.89, 700, 2500
Resolution, mm3 1.0×1.0×1.0 1.0×1.0×1.2
Average lesion intensity Q2 (IQR) 1006 (879–1130) 1517 (1256–1758)
3.2 Model
We chose nnU-Net, a widely adopted framework for medical image segmentation [27, 28], previously explored for lesion segmentation specifically [24, 44]. Based on the best practices for medical image segmentation, nnU-Net automatically configures pre-processing, training hyperparameters, and post-processing for a specific dataset. This approach ensures state-of-the-art results without extensive manual intervention, yet it allows for modifications on demand. We used the standard nnU-Net architecture operating on 3D full-resolution MP2RAGE scans.
Given small sizes of CLs and, thus, high class imbalance, the convergence of nnU-Net benefited from adjusting the optimization procedure. The AdamW optimizer was chosen for its ability to handle sparse gradients and class imbalances, with decoupled weight decay providing improved optimization stability for challenging datasets. The learning rate was reduced to 0.0001 to ensure stable convergence and enhance segmentation accuracy, particularly for minority classes and anatomically challenging regions. We provide the implementations in our GitHub: https://github.com/NataliiaMolch/interpret-lesion-unc.
3.2.1 Performance evaluation
Several measures chosen based on previous works were used to evaluate the model. These measures compare the predicted and ground-truth segmentation: (normalized) Dice similarity score (DSC and nDSC), lesion detection F1-score (LF1), precision (LPPV), and recall (LTPR). Analogously to DSC, nDSC measures the overall segmentation quality, however, it corrects for the difference in lesion loads across subjects, making their comparison fair [61]. For the lesion detection measures, the lesion regions were defined using the connected component analysis with 26 connectivity. True positive lesions were defined as lesions with a non-zero overlap with the ground truth. False positive lesions are predicted connected components without an overlap with the ground truth. False negative lesions are ground truth lesions without an overlap with the predicted lesion mask. The detailed metrics definition is described in A.
3.3 Uncertainty quantification
3.3.1 Deep ensemble
Deep ensembles for UQ have a wide adoption for medical imaging tasks given a strong uncertainty-robustness link and higher calibration compared to other methods, like temperature scaling, test time augmentation, Monte Carlo dropout, or stochastic weight averaging [8, 4]. Deep ensembles are a non-Bayesian UQ method based on training several models with different random seeds. Thus, each model in the ensemble has a distinct weight initialization, augmentations, training examples, shuffling, etc., providing convergence to different local minima. The resulting members’ weights are treated as sample estimations from the posterior distribution of model parameters. For each image voxel, uncertainty is computed as a ”spread” of the members’ predictions given the same example, i.e., using information theory measures (e.g., variance, mutual information, etc). We implemented UQ using a deep ensemble of $K=5$ models based on the nnU-Net framework architecture, training each model independently without using the bagging technique.
3.3.2 Lesion-scale uncertainty measures
Semantic segmentation task models output a tensor, associating class probabilities with each pixel or voxel of the input image. Such spatial nature allows uncertainty to be quantified at multiple levels: per voxel or ROI (e.g., focal lesions). While pixel-wise uncertainty provides detailed spatial information, instance-level uncertainty (computing a single uncertainty score for each predicted lesion) offers more clinically actionable insights. Computing a single uncertainty value per predicted lesion condenses predictive variability into compact, clinically meaningful information that aligns with how radiologists interpret and report findings, focusing on individual lesions rather than isolated voxels [13].
We quantify the uncertainty associated with each predicted CL using the lesion structural uncertainty (LSU) measure, which we previously proposed for WML segmentation [61]. LSU showed a slightly better ability to capture lesion false discovery errors. Additionally, having a simple interpretable definition and being bounded from 0 to 1, LSU allows for comparison across different patients. LSU quantifies the average disagreement between each ensemble member (k) lesion prediction, $L^{k},k\in\{0,2,\dots,K-1\}$, and the ensemble’s aggregated prediction $L$. Here, $L$ represents the final lesion segmentation obtained by averaging the probability maps from all $K$ ensemble members and applying a threshold to obtain a binary segmentation mask. LSU is defined as follows:
$LSU=1-\frac{1}{K}\sum\limits_{k=0}^{K-1}IoU(L,L^{k}),$
where $IoU(\cdot,\cdot)$ denotes the intersection over the union.
3.4 Uncertainty explainability analysis framework
We aimed to explain lesion-scale uncertainty values, described by LSU, by quantifying their relationship with diverse lesion and patient characteristics. We selected lesion features that reflect best practices in medical image region characterization, enriched by clinically relevant metrics. These features were used to explain the variability of uncertainty using a linear or random forest model. Feature importance analysis brings insights into the information carried by predictive uncertainty. We investigated the variability of the proposed analysis in different settings: training data domain, in-domain testing data, and out-of-domain testing data. We further examined the relationship between the important features and the underlying sources of doubt in expert rater decisions. Figure 2 illustrates the proposed framework. The code is provided in the aforementioned GitHub.
Figure 2: Illustration of the proposed framework for explaining instance-wise uncertainty of a DL segmentation model.
3.4.1 Feature engineering
Each predicted CL was characterized by its segmentation quality, location, intensity, texture, shape, surrounding intensity, and degree of novelty.
Segmentation quality was computed as an adjusted intersection over the union (IoUadj[62] between the predicted lesion region ($L$) and the ground truth lesion mask. IoUadj is a suitable measure for medical image segmentation, where multiple ground truth instances can overlap with multiple predicted instances. IoUadj corrects the union computation, removing the ground truth part explained by other predicted instances. The IoUadj computation is detailed in B.
The lesion location was quantified using the MNI atlas [63], separated into the right and left hemispheres. The location features were computed as the intersection between the lesion region and the MNI brain region divided by the lesion size in voxels. The location with respect to the cortex (i.e., cortex involvement) was computed as the intersection normalized by lesion size with the GM defined by SynthSeg [64]. In radiological practice, the unclear cortical involvement contributes to the aleatoric uncertainty in distinguishing between CL and WML that only touch the cortex (i.e., juxtocortical lesions).
Radiomics features suggested by the PyRadiomics framework computes intensity and shape features for lesions and their surroundings [65]. Lesion surrounding was defined as a 4-voxel-wide region around the lesion, which was chosen in agreement with the medical doctors with an admitted degree of ambiguity. We used a morphological dilation with a $3\times 3\times 3$ voxels to obtain the lesion surrounding. PyRadiomics intensity features (first-order statistics) describe the distribution of voxel intensities within the ROI, such as mean, variance, skewness, and energy, providing insights into the overall brightness, variability, and histogram shape. The texture features (second-order statistics) analyze the spatial relationship and patterns of voxel intensities at specific distances and directions. We focused on the texture features derived from the gray-level co-occurrence matrix (GLCM) with a distance of 1 voxel in all 13 3D directions, quantifying contrast, correlation, energy, and homogeneity. The rest of the texture features are omitted given the low interpretability and to limit the amount of features below the number of predicted lesions. The shape features capture the 3D geometry of the ROI and focus on the structural and morphological aspects, e.g., volume, surface area, fraction (compactness, sphericity), and longest or shortest axes, among others.
The degree of novelty describes the difference between a lesion and lesions in the Train set. To measure the lesion novelty, we computed the distance between the lesion and the training set using the engineered lesion and perilesional features. For dimensionality and noise reduction, probabilistic principal component analysis (PPCA) [66] was applied to the features before computing the distance. PPCA provides a maximum likelihood estimation approach to dimensionality reduction while accounting for noise in the data. We used two measures to compute the distance: Mahalanobis distance and non-zero distance to the closest example in the training set.
Patient-wise measures are also used to examine how global features are related to the lesion uncertainty. First, we added PyRadiomics intensity statistics for the whole predicted lesion region. Second, to study the relationship between uncertainty and image quality, we relied on MRIQC [67], characterizing various aspects of MR image quality such as signal-to-noise ratio, motion artifacts, ghosting, and entropy in the whole brain and the specific brain tissues. We selected 13 metrics applicable for structural skull-stripped MRI scans (the selection procedure is detailed in the Supplementary materials). Finally, we used patient demographic and clinical information, such as sex, age, and expanded disability status scale (EDSS).
The total number of selected features was N=156. Table 2 summarizes all the features.
Table 2: Computed features studied in relationship with the lesion-scale uncertainty.
Type Number of features ROI Description Examples
Prediction quality 1 lesion Adjusted intersection over the union between the predicted lesion and the ground truth lesion mask IoUadj
Location 19 lesion Intersection between the lesion and atlas-defined brain regions divided by the lesion size GM overlap, temporal/frontal/occipital left or right lobes overlap
Intensity 18 × 3 lesion, perilesional region, predicted CL mask Characterize the intensity distribution within ROI Mean, standard deviation, skewness, percentiles, energy
Texture 24 × 2 lesion, perilesional region Characterize the distribution of co-occurring voxel values within ROI Contrast, cluster prominence, entropy
Shape 14 lesion Structural and morphological aspects of ROI’s 3D geometry Volume, surface area, volume-to-surface ratio, flatness, sphericity, maximum diameter
Novelty 2 lesion Distance to the training set lesions Mahalanobis distance or distance to the closest training example
Image quality 13 brain, GM, WM MRI quality control measures Signal-to-noise ratio in brain and brain tissues, intensity non-uniformity (INU) median, entropy focus criterion (EFC)
Demographics 5 patient Clinically relevant features Age, male sex, female sex, MS duration, EDSS score
3.4.2 Feature processing
The computed feature table contained missing values in the Test-in and Test-out sets. The missing values appeared due to the absence of demographic information (EDSS score missing for 1 patient in Test-in and the second timepoint within Test-out; disease duration missing for 3 patients in Test-out) or due to PyRadiomics not producing the results (1 lesion in Test-in and 2 lesions in Test-out). Missing values were imputed using a k-nearest-neighbors approach with five neighbors and uniform weights. The imputation was done before fitting the uncertainty regression model. While imputation before model fitting would typically risk data leakage in predictive tasks, such an approach is beneficial for explainability purposes, providing more reliable estimates of missing values. Before fitting an explainer model, features with low variance were filtered out. The $10^{-6}$ variance threshold was used based on the 32-bit float number precision. The features were standardized separately for each dataset, which makes the coefficients of the linear model comparable to each other.
3.4.3 Uncertainty regression model
We explored both linear regression and random forest models to study the relationship between uncertainty (dependent variable) and relevant lesion/patient features (independent variables). On one hand, linear regression provides a simple and interpretable framework with clear insights into the relationship between variables. On the other hand, random forests can model complex and non-linear relationships between independent and dependent variables, albeit at the expense of explainability.
Linear model. We chose Elastic-Net [68] model, which combines L1 and L2 regularization, providing feature selection and penalizing for multi-collinearity, respectively. Given the higher prevalence of low-uncertainty lesions and a greater interest in explaining high-uncertainty lesions, we implemented an increased penalty for high-uncertainty examples, i.e., sample weighing based on uncertainty in the loss function and the quality metrics. The model selection procedure was performed using grid search, optimizing the weighted $R^{2}$ score obtained on 5-fold cross-validation. The hyperparameters tuned during the model selection control the regularization strength (9 values from $10^{-4}$ to $10^{2}$ evenly spaced on the log scale), the L1 to L2 regularization strength ratio (9 values evenly spaced between 0 and 1), and the tolerance for the optimization ($10^{-3}$, $10^{-4}$, $10^{-5}$). To obtain robust feature importance values, the grid search procedure was repeated 10 times with different random seed initializations, controlling the data-splitting and Elastic-Net convergence. Additionally, to study the relationship between the prediction quality and uncertainty, we used the ordinary least squares regression model. The analysis was repeated 10 times with the random seed change affecting the cross-validation data-splitting.
Random forests. Random forests [69] is an ensemble learning method that builds multiple decision trees during training, using random subsets of data and features, and combines their outputs (via averaging for regression or majority voting for classification) to make predictions. Feature importances in random forests are computed by evaluating how much each feature reduces the impurity (in our regression case, variance) across all trees in the forest. The model selection procedure was performed using grid search, optimizing the $R^{2}$ score obtained on 5-fold cross-validation. The search grid was defined by varying the number of estimators (20, 50, 100), the maximum depth of the trees (none, 5, 10), the minimum number of samples required to split an internal node or to be at a leaf node (2, 5), and the number of considered features for splitting (N, N²). Analogously to the linear model, the grid search was repeated 10 times with different random seed initialization.
3.4.4 Feature importance
The coefficients of the linear models were used to evaluate the strength of the relationship between features and uncertainty, given that the features have a similar scale. For each of the datasets, i.e., Train, Test-in, Test-out, the coefficients were obtained during each of 10 runs, corresponding to different random seeds. During each run, the model selection procedure was performed (grid search with 5-fold cross-validation $R^{2}$ optimization) and the model was fitted on the dataset. Reported feature importances were obtained by averaging across 10 runs, the standard deviation was computed for each selected feature to evaluate the variability.
3.5 Evaluation of explanations
3.5.1 Quality of fit
To assess the quality of the model fitting during the model selection, the coefficient of determination $R^{2}$ was used, which represents the proportion of the variance in the dependent variable (i.e. LSU) that is explained by the independent variables (i.e. lesion or patient features) in a linear regression model:
$R^{2}=1-\frac{SS_{res}}{SS_{tot}}=1-\frac{\sum_{i=0}^{n-1}w_{i}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=0}^{n-1}w_{i}(y_{i}-\bar{y}_{i})^{2}},$
where $y_{i}$, $\hat{y}_{i}$ are true and predicted values, $SS_{res}$ is the residual sum of squares or unexplained variance, $SS_{tot}$ is the total sum of squares or total variance in the observed variable, $\bar{y}=\frac{1}{n}\sum_{i=0}^{n-1}w_{i}y_{i}$, $w_{i}$ sample weights. If $R^{2}$ is unweighed $w_{i}=1,\quad i=0,\dots,n-1$.
$R^{2}=1$ indicates the best possible quality of fit, and $R^{2}<0$ indicates that the model is worse than using the mean of the dependent variable y as a prediction. Additionally, the error magnitude was reported using the mean absolute error (MAE):
$MAE=\frac{1}{n}\sum_{i=1}^{n}w_{i}|y_{i}-\hat{y}_{i}|$
The quality of fit as a proxy to the quality of explanations was computed for each of 10 runs, either during the 5-fold cross-validation or over a dataset (Train, Test-in, Test-out) for the best selected model. We reported the average and standard deviation across 10 runs with different random seeds.
3.5.2 Expert rater feedback
We used feedback from the experts annotating CLs to study the relationship between the factors underlying model uncertainty and the challenges in manual CL annotation. We aimed to identify the factors that may hinder medical doctors’ confidence in annotating CLs. We assume that any factor should lie in one of five categories: i) intensity and contrast, ii) location in the brain, iii) location with respect to the cortex, iv) lesion shape, and v) patient-related information. Using the domain knowledge, available guidelines, our prior research, and unstructured interviews with doctors, we proposed several factors for each category, leaving the possibility for the expert raters to suggest other factors. For each of the categories, the respondents needed to select all the factors that decrease either confidence in deciding if an object is a CL or the overall visibility of CLs, making them difficult to segment or detect. We left to the respondents the right to reject any of the factors or propose new ones. All the questions can be found in Figure 3. In total, five medical doctors with different levels of expertise underwent the questionnaire: one neuroscientist with 3 years of experience and 4 neurologists with 6, 9, 10, and 20 years of experience in MS.
Figure 3: The clinical questionnaire with guidelines provided to the doctors on the left and questions on the right.
4 Results
4.1 Deep learning segmentation model performance
Table 3 reports CL segmentation and detection performance of the trained nnU-Net model. Judging by the Dice similarity score (DSC) and detection F1, the performance on Test-out was significantly worse than on Test-in. This confirms the suitability of Test-out for investigating the quality of uncertainty out-of-domain, since it enables the evaluation of uncertainty explanations and sanity in the scenario of model failure. The total number of predicted lesions is also reported as the number of examples used to fit the explainer model.
Table 3: Mean and standard deviation of the segmentation (DSC, nDSC) and lesion detection quality metrics (LF1, LPPV, LTPR) across Train, Test-in, and Test-out datasets, as well as the number of predicted CL (# CLs).
Dataset DSC nDSC LF1 LPPV LTPR # CLs
Train 0.908 $\pm$ 0.082 0.740 $\pm$ 0.289 0.957 $\pm$ 0.078 0.991 $\pm$ 0.030 0.934 $\pm$ 0.120 753
Test-in 0.617 $\pm$ 0.332 0.494 $\pm$ 0.380 0.728 $\pm$ 0.300 0.759 $\pm$ 0.333 0.778 $\pm$ 0.287 284
Test-out 0.318 $\pm$ 0.373 0.284 $\pm$ 0.372 0.399 $\pm$ 0.396 0.529 $\pm$ 0.464 0.377 $\pm$ 0.406 273
4.2 Performance in explaining uncertainty
The quality of fit is presented in Table 4 for linear models. The contribution of different feature subgroups was investigated, including i) all features; ii) only prediction quality measured by IoUadj; iii) only lesion-scale features, i.e., all the patient-scale features were excluded; and iv) all features excluding prediction quality.
For the linear model, using all features, the $R^{2}$ was highest when fitted on the Train set; it decreased by 20% on Test-in and by an additional 33% on Test-out. The MAE increased by 50% from Train to Test-in and by 17% from Test-in to Test-out. The standard deviation of $R^{2}$ on the Test-out set is high due to an outlier appearing under a specific random seed. However, this outlier did not affect the feature importance analysis (see Supplementary materials).
The quality of fit, measured by $R^{2}$, more than doubled when all features were used compared to using only lesion segmentation quality to predict uncertainty. This suggests that the designed features captured more variability in uncertainty than lesion segmentation quality (IoUadj) alone. Fitting the explainer on all features, but the predicted quality led to up to an 11% decrease in the $R^{2}$ compared to using all features on Test-in and Test-out. This indicates that the proposed features could not completely replace the information about the segmentation error. When patient scale information was omitted, $R^{2}$ improved by 2% on Train and 3% on Test-out, but decreased by 2% on Test-in.
Table 4: The quality of fit of the linear models measured by the coefficient of determination $R^{2}$ ($\uparrow$) and $MAE$ ($\downarrow$) computed on cross-validation on Train, Test-in, and Test-out sets. The mean and standard deviation are computed across 10 runs with different random seeds. Different feature spaces are explored, including all features, only lesion-scale features, and all features without prediction quality. Best values per column are highlighted in bold.
Feature set Train Test-in Test-out
$\mathbf{R^{2}}$ MAE $\mathbf{R^{2}}$ MAE $\mathbf{R^{2}}$ MAE
All 0.571 $\pm$ 0.039 0.088 $\pm$ 0.001 0.465 $\pm$ 0.046 0.132 $\pm$ 0.003 0.309 $\pm$ 0.202 0.155 $\pm$ 0.005
Only IoUadj 0.237 $\pm$ 0.062 0.125 $\pm$ 0.001 0.227 $\pm$ 0.067 0.168 $\pm$ 0.003 -0.012 $\pm$ 0.252 0.192 $\pm$ 0.005
Only lesion-scale 0.580 $\pm$ 0.039 0.087 $\pm$ 0.001 0.458 $\pm$ 0.043 0.137 $\pm$ 0.004 0.318 $\pm$ 0.181 0.153 $\pm$ 0.005
No IoUadj 0.568 $\pm$ 0.037 0.089 $\pm$ 0.001 0.414 $\pm$ 0.044 0.139 $\pm$ 0.005 0.278 $\pm$ 0.249 0.160 $\pm$ 0.006
The random forest quality of fit is shown in Table 5. MAE values of random forest models are comparable with the values for the linear models for Train and Test-in. The $R^{2}$ is lower compared to the linear model for each feature set: up to 15% lower on Train and Test-in, up to 60% lower on Test-out. For all datasets, removing either patient-scale information or the prediction quality led to lower $R^{2}$ values compared to using all features; however, the difference was marginally small for Train and Test-in. Thus, further, we reserved the linear model for explanations in the setting with all features. Despite slightly lower performance, the standard deviation of the cross-validation scores across different seeds is lower for the random forest model compared to linear models.
Table 5: The quality of fit of the random forest models measured by the coefficient of determination $R^{2}$ (↑) and $MAE$ (↓) computed on cross-validation on Train, Test-in, and Test-out sets. The mean and standard deviation are computed across different random seeds. Different feature spaces are explored, including all features, only lesion-scale features, and all features without prediction quality. Best values per column are highlighted in bold.
Feature set Train Test-in Test-out
$\mathbf{R^{2}}$ MAE $\mathbf{R^{2}}$ MAE $\mathbf{R^{2}}$ MAE
All 0.504 $\pm$ 0.031 0.089 $\pm$ 0.003 0.398 $\pm$ 0.036 0.135 $\pm$ 0.005 0.184 $\pm$ 0.084 0.167 $\pm$ 0.006
Only lesion-scale 0.497 $\pm$ 0.025 0.090 $\pm$ 0.002 0.391 $\pm$ 0.040 0.136 $\pm$ 0.005 0.121 $\pm$ 0.084 0.172 $\pm$ 0.008
No IoUadj 0.495 $\pm$ 0.025 0.090 $\pm$ 0.003 0.356 $\pm$ 0.035 0.141 $\pm$ 0.003 0.181 $\pm$ 0.086 0.168 $\pm$ 0.007
We analyze the linear model prediction quality on sets different from the fitting sets and evaluate the quality of explanations under the domain shift. The results are shown in Figure 4. A drop in performance caused by the domain shift was observed for all fitting settings, except for one. When IoUadj was used to predict uncertainty using Test-out for fitting, the $R^{2}$ was higher on Train and Test-in compared to Test-out. This should indicate a weaker relationship between uncertainty and error on the out-of-domain set, compared to in-domain sets. The drop in performance tended to be worsened by the out-of-domain setting. Thus, explanations on the Train and Test-in sets should be more similar to each other than to Test-out.
Figure 4: The quality of fit of the linear models measured by the coefficient of determination $R^{2}$ ($\uparrow$, red) and MAE ($\downarrow$, blue) computed on Train, Test-in, and Test-out sets. Rows correspond to different pairs of sets used for fitting a linear model and different feature sets; columns correspond to the sets used for the evaluation of the linear model. The mean and standard deviation are computed across different random seeds.
4.3 Feature importance analysis
Figure 5 presents coefficients of the Elastic-Net model fitted to the entire feature space, ranked by magnitude; 20 coefficients explaining the biggest part of variance are displayed. Since all the features were standardized before model fitting, the absolute values are comparable and indicate feature importance. The features with the highest absolute importance explain the most variability in uncertainty. Positive values coefficients indicate features that correlate with increased uncertainty (when the feature value rises, uncertainty typically increases), while negative coefficients represent features associated with decreased uncertainty.
Figure 5: Barplot with linear explainer coefficients treated as feature importance. Barplots are built for Train, Test-in, and Test-out sets. Positive values - direct relationship with uncertainty, and negative values - direct relationship with certainty. Twenty features with the highest absolute coefficients are built.
There were similarities in the selected features between different fitting sets (Train, Test-in, and Test-out). The surface-to-volume ratio explained the variability in uncertainty across all fitting sets and corresponded to higher uncertainty values. The surface-to-volume ratio describes both the shape and size of the object. High values of this feature correspond to irregular or elongated shapes and smaller sizes of the objects. Other features describing shape and size emerged as relevant predictors of lesion uncertainty for the different datasets. Especially on the test sets, more spherical lesions (high flatness and elongation PyRadiomics features and low surface-to-volume ratio) of bigger sizes (high maximum 2D diameter and minor axis length, low surface-to-volume ratio) were related to higher certainty. Prediction quality measured by IoUadj was another high-importance feature shared by all the fitting sets, with a higher IoUadj value corresponding to lower lesion uncertainty. GM overlap appears in the analysis for all sets. Higher values of GM overlap coincided with lower uncertainty. Lesion novelty feature (smallest distance) appeared for both Train and Test-in sets and had a mediocre importance, indicating that lesions dissimilar to the Train set lesions have higher uncertainty. Different intensity features were selected during the feature importance analysis. They are related not only to the lesion region itself, but also to the lesion surrounding, and the whole predicted lesion region. The particular features selected vary between different sets. Generally, the presence of (outlying) high intensity within a lesion indicated lower uncertainty (high lesion intensity 90% percentile, variance, mean absolute deviation, maximum, range). Perilesional intensity kurtosis was selected for all fitting sets, however, it had a low relative importance. Higher perilesional intensity kurtosis was associated with lower uncertainty values. Some feature groups were not selected or have relatively low importance, including texture, location in the brain, patient characteristics, and image quality (except for the Test-in set). While Test-in and Test-out had some similarities in the selected features, there were some visible differences. First, a weaker relationship between uncertainty and error in Test-out, judging by the lower relative importance of IoUadj. Additionally, for Test-out, the accumulated importance of the perilesional intensity features was higher than for shape features or IoUadj. The feature importance plots for the random forest model are provided in the Supplementary materials.
4.4 Selected features analysis
The joint distribution of uncertainty and high-importance features in different datasets is shown in Figure 6. The median Pearson correlation between uncertainty and surface-to-volume ratio is 0.61, and IoUadj is 0.54. Compared to the in-domain sets, the correlation with IoUadj is 35% lower on the out-of-domain Test-out. Thus, the relationship between uncertainty and lesion prediction quality, measured by IoUadj, is weaker on the out-of-domain set. GM overlap had a weak correlation with lesion uncertainty (r=-0.07 for Train, -0.21 for Test-in, and -0.04 for Test-out). The shape of the distribution indicates that the relationship is not linear. Lesion novelty measured by the Smallest distance feature had a weak positive correlation ranging from 0.13 to 0.32 Pearson correlation coefficient.
Figure 6: Bivariate histograms with joint distributions of features and lesion uncertainty for surface-to-volume ratio, IoUadj, GM overlap, and (lesion novelty) smallest distance features computed on Train, Test-in, and Test-out data. Pearson correlation $r$ is computed on the standardized features.
To better understand the information contained in features, we used lesion visualization. Figure 7 shows lesion examples obtained based on the feature distributions for shape, location, and intensity high-importance features. The feature with the highest importance, i.e., surface-to-volume ratio, clearly communicates the lesion size; lesion elongation could also be captured by this feature. The GM overlap feature, important for all sets, is reflective of the juxtacortical (a WML, adjacent to the cortex, but contained in WM), leukocortical (a CL, overlapping with GM and WM), and intracortical lesions (contained in GM). This said, the segmentation of GM is subject to errors and partial volume effect, potentially impairing the reliability of the derivative feature. We find that all the highest-importance intensity features carried some information about the location with respect to GM. The perilesional intensity median indicated a greater border with WM while being associated with higher uncertainty. Thus, this feature could be a proxy for the juxtacortical versus CL distinction. Similarly, the mean absolute deviation of the lesion intensity and perilesional intensity IQR possibly served this distinction.
Figure 7: Lesion examples sampled based on the feature distribution. The shape, location, and intensity features were selected based on the feature importance analysis. CLs were obtained based on 0, 25, 50, 75, and 100 percentiles (columns from left to right) of each feature distribution on a particular set (rows - Train, Test-in, or Test-out). All the features have a positive relationship with lesion-wise uncertainty, except for GM overlap and lesion intensity mean absolute deviation, which is negatively related to uncertainty. Only lesions with a non-zero overlap with the ground truth are displayed.
4.5 Expert rater questionnaire
Figure 8 summarizes the results of the expert rater questionnaire, identifying key factors contributing to the low visibility and reduced manual segmentation quality of CLs on MP2RAGE. The proposed concepts, were additioned by the respondents with respect to the ”Location in the brain” category. The proposed concepts include i) similarity in appearance to perivascular spaces in the insular cortex and ii) temporal lobes.
The responses indicated that subpial lesions, small lesions, lesions with low contrast with the surrounding tissue, and scans with low resolution are the most challenging to annotate (4 out of 5 raters). Lesion location relative to the cortex was also selected by the majority of participants, suggesting that both lesions with unclear cortical involvement and intracortical lesions pose significant challenges for annotators. Image quality-related factors such as artifacts and blurriness were also selected by the majority of respondents as additional obstacles. Furthermore, the presence of high-intensity values or heterogeneous intensity patterns within a lesion was noted. Two experts specifically highlighted the insular cortex and temporal lobes as brain regions where CLs are more difficult to annotate. None of the respondents selected patient-related factors as influencing annotation difficulty, and all reported that patient-specific information was not accessed before annotation.
Figure 8: Barplot summarizing the expert rater feedback completed by five MP2RAGE CL annotators. The different colors correspond to different categories of features from the questionnaire.
5 Discussion
The proposed framework, which regresses uncertainty from human-interpretable features, explores the information encoded by instance-wise deep ensemble uncertainty for multiple sclerosis cortical lesion segmentation on MRI.
5.1 Information carried by instance-wise uncertainty
Our findings demonstrated that deep ensemble uncertainty is not solely explained by the quality of the prediction (IoUadj) or distance from the training set (i.e., lesion novelty). Instead, various features characterizing the size, shape (surface-to-volume ratio), and location of the lesion with respect to the cortex (GM overlap and various intensity features) were more important in explaining the variability of uncertainty. The selected features align with the main factors that hinder the confidence of the annotators, discovered with the feedback from medical experts: small and elongated (subpial) lesions and the distinction between juxtacortical and CL (i.e., cortical involvement).
This alignment between model uncertainty and annotator uncertainty suggests that deep ensemble uncertainty also encodes aleatoric noise. Deep ensemble uncertainty is commonly associated with epistemic uncertainty and, thus, more strongly correlated with model error than other UQ methods [37, 70]. However, this assumptions stems largely from low-noise tasks. In contrast, CL segmentation is inherently noisy. Moreover, hidden data biases, such as underrepresentation of certain lesion subtypes (e.g., subpial or small intracortical lesions) [30, 71], may also contribute to increased uncertainty. This interplay of factors raises questions about the separability of the sources of aleatoric and epistemic uncertainty [72].
As discussed in Section 4.4, non-linear dependencies are present between uncertainty and selected features with high importance. Partially, an unexplained uncertainty should be attributed to the non-linear relationship. However, modelling these dependencies requires more data, since random forests explainers did not outperform linear models in goodness of fit, likely due to overfitting in small datasets [73]. While linear models are pragmatic for limited data, their results require supplementary analysis (e.g., partial dependence plots) to characterize non-linear dependencies.
5.2 Explainability of uncertainty under domain shifts
We evaluated uncertainty and the proposed framework performance under domain shift — a clinically realistic scenario where prediction quality varies across settings. As shown in Section 4.2, the explanatory power of uncertainty degraded progressively from Train to Test-in to Test-out, indicating reduced informativeness with growing domain shift. We hypothesize that this degradation is caused by changes in relevant features and a deterioration in uncertainty quality. Prior work already showed that UQ reliability declines under distributional shifts [70, 54, 48].
Our results further suggest that uncertainty estimates under strong domain shift may lose informativeness about both prediction error and aleatoric noise, given higher unexplained uncertainty on Test-out (Tables 4 and 5). We also observed performance degradation when an explainer was fitted to one set and applied to another. The drop in performance quantified the transferability of explanations between Train and Test-in sets (i.e., in-domain sets), which is better than the in-domain versus Test-out. These findings suggest that uncertainty explainability is sensitive to domain shift, both in terms of uncertainty quality and explainer reliability.
5.3 Predictive uncertainty and annotator confidence
We found a strong correspondence between the features associated with high predictive uncertainty (Section 4.3) and the factors underlying low annotator confidence (Section 3.5.2). An increased surface-to-volume ratio characterizes small lesions and subpial lesions (elongated, extending along the cortex). As shown in Sections 4.3, the surface-to-volume ratio feature was associated with higher lesion uncertainty. GM overlap (i.e., cortical involvement) and intensity-based features were relevant for both annotator confidence and model prediction certainty.
However, a direct one-to-one mapping between clinically identified factors and engineered lesion features is not feasible. The importance of lesion location, as emphasized by experienced raters, also extends to predictive uncertainty. The non-linear relationship between lesion uncertainty and GM overlap could be explained by the fact that the model is both uncertain on lesions with unclear cortex involvement (GM overlap 0) and on the intracortical lesions (GM overlap = 1). Translating engineering characteristics into understandable concepts is not straightforward [74] and beyond the scope of this work.
5.4 Future implications of uncertainty explanations
Explainability of uncertainty has practical values in many different domains. From an engineering perspective, characterizing difficult and ambiguous cases may contribute to the development of methods to improve model robustness or uncertainty calibration. From the stakeholder perspective, it supports human-centered AI development by enabling transparency, fostering trust, and integrating expert feedback.
On the one hand, explaining uncertainty and visualizing important features fosters communication with medical professionals. On the other hand, expert feedback enriches our understanding of model behavior, revealing its data-dependent nature.
Our results suggests that unexplained uncertainty is partially attributed to the approximation error. This means that some uncertainty values are misleading and have the power to bias medical decision-making into the wrong direction. Quantifying this unexplained component could support trust calibration, aligning human trust with actual model reliability. However, further work is needed to evaluate how clinicians perceive and utilize such explanations.
While our framework was demonstrated in the context of MRI-based lesion segmentation, it is easily adaptable to other UQ methods, machine learning tasks (e.g., classification), imaging modalities, and clinical applications.
Finally, while the proposed analysis is correlation-based, it makes hypotheses about the causal relationship, which should be further investigated in future work.
6 Conclusion
We proposed a novel framework for explaining AI-models predictive uncertainty, beyond classical error estimation analysis. Our results on lesion uncertainty confirm the relationship between uncertainty and error and link high uncertainty with clinically relevant concepts. We also show the limitation of UQ in the out-of-domain setting and assess the presence of misleading uncertainty values. Overall, increasing the understanding of AI model reliability in medical image analysis is needed to increase transparency, facilitate the human-AI interactions, and ultimately support building stakeholders trust in AI decisions.
Acknowledgment
This work was supported by the Hasler Foundation Responsible AI program (MSxplain) and the Research Commission of the Faculty of Biology and Medicine (CRFBM) of UNIL. This research was partially supported by the Intramural Program of NINDS, NIH, USA. We acknowledge access to the facilities and expertise of the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by Lausanne University Hospital (CHUV), University of Lausanne (UNIL), École polytechnique fédérale de Lausanne (EPFL), University of Geneva (UNIGE) and Geneva University Hospitals (HUG). The authors wish to thank Dr. Mara Graziani for valuable consultations and discussions at the initial stage of this research.
Appendix A Metrics for model performance evaluation
Let TP, FP, FN be the number of true positives, false positives, and false negatives. The quality metrics are defined as follows:
$F_{1}=\frac{2TP}{2TP+FP+FN},$
$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN}.$
In the case of segmentation, the F1 score is called the Dice similarity score; in case of lesion detection, the F1, precision, and recall are called, LF1, LPPV, and LTPR, respectively.
Lesion detection metrics rely on TP, FP, and FN lesion definitions. For the lesion detection measures, the lesion regions were defined using the connected component analysis with 26 connectivity. True positive lesions were defined as lesions with a non-zero overlap with the ground truth. False positive lesions are predicted connected components without an overlap with the ground truth. False negative lesions are ground truth lesions without an overlap with the predicted lesion mask.
Normalized Dice similarity score proposed by Raina et al. (2023):
$F_{1}=\frac{2TP}{2TP+\kappa\cdot FP+FN},\quad\kappa=h(r^{-1}-1)$
where $h$ is the ratio between the positive and the negative classes, $r\in(0,1)$ is the reference value set to the mean fraction of the positive class.
In our case, it is the lesion load: $r=2\cdot 10^{-5}.$
Appendix B Adjusted intersection over the union definition
Let $x$ be an input image and let $\hat{S}_{x}=\{\hat{y}_{z}(x,w)\mid z\in x\}$ be the pixel-wise predicted segmentation, where $\hat{y}_{z}(x,w)$ is the predicted class at pixel $z$ given model parameters $w$. Define:
  • 1.
    $\hat{K}_{x}$ — the set of connected components (segments) in the predicted segmentation $\hat{S}_{x}$
  • 2.
    $K_{x}$ — the set of connected components in the ground truth segmentation $S_{x}$
For a predicted segment $k\in\hat{K}_{x}$, define:
  • 1.
    $K^{\prime}=\bigcup\{k^{\prime}\in K_{x}\mid k^{\prime}\cap k\neq\emptyset\text{ and }\text{class}(k^{\prime})=\text{class}(k)\}$ — union of all ground truth components that overlap with $k$ and have the same class
  • 2.
    $Q=\{q\in\hat{K}_{x}\mid q\cap K^{\prime}\neq\emptyset\}$ — set of all predicted segments that overlap with $K^{\prime}$
Then, the adjusted intersection over union is defined as:
$\mathrm{IoU_{adj}}(k)=\frac{|k\cap K^{\prime}|}{|k\cup(K^{\prime}\setminus Q)|}$
This formulation corrects the standard intersection over union by not penalizing predicted segments for missing parts of the ground truth that are already covered by other predicted segments of the same class.
Ethics statement
Studies involving human data were approved by the local ethics committees; informed consent was obtained from all participants before study entry.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Grammarly and GhatGPT for grammatical and stylistical errors correction. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.